OpenStudy (anonymous):

giving medals!!!!!! write an equation of the tangent function with period 3pi/8, phase shift -pi/5, and vertical shift -2 a. y=tan(8x/3-8pi/15)-2 b. y=tan(8x/3+8pi/15)-2 c. y=tan(16x/3+3pi/80)-2 d. y=tan(8x/3-3pi/40)-2

OpenStudy (phi):

equation of tan function: y=tan(Bx-C), period=pi/B, phase shift=C/B